5. Suppose you have the following production function: Q = f (L, K) = 10L ½ K ½ . In addition, the price of labor is $1 and the price of capital is $4

a. What is the optimal amount of labor and capital if you want to produce 20 units?

b. What is the level of minimum cost ?( Ans L=4 and K=1,Min C=$8)

6. Suppose the short run production function can be represented by Q = 60,000L 2 – 1000L 3 . Then, determine

a. The level labor employment that maximizes the level of output

b. The level of employment that maximizes APL and the maximum APL

1
Expert's answer
2021-12-22T11:54:40-0500

Solution:

a.). Cost minimization:

MRTS = "\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


TC = 10L ½ K½


MPL = "\frac{\partial Q} {\partial L}" = 5L-0.5 K0.5


MPK = "\frac{\partial Q} {\partial K}" = 5L0.5 K-0.5


"\frac{MP_{L} }{MP_{K}} = \frac{w }{r}"


w = 20

r = 80


"\frac{5L^{-0.5} K^{0.5} }{5L^{0.5} K^{-0.5}} = \frac{20 }{80}"


"\frac{K }{L} = 0.25"


K = 0.25L

Q = 10L ½ K½

20 = 10(L0.5) (0.25L0.5)


L = 8

K = 0.25L = 0.25(8) = 2

The optimal amount of labor and capital to produce 2000 units (L,K) = (8, 2)

Labor = 8

Capital = 2


b.). Minimum cost:

C = wL + rK

C = (20 "\times" 8) + (80 "\times" 2) = 160 + 160 = 320

Minimum cost = 320


b.). The level of employment that maximizes APL and the maximum APL:

APL = "\frac{Q}{L}" = 60,000L – 1,000L2

Maximize APL:


"\frac{\partial AP_{L} } {\partial L}" = 60,000 – 2,000L

Set APL = 0

60,000 – 2,000L = 0

60,000 = 2,000L

L = 30

The level of employment that maximizes APL = 30 units


APL = 60,000L – 1,000L2 = 60,000(30) – 1000(302) = 1,800,000 – 900,000 = 900,000

APL = 900,000